OpenStudy (anonymous):

6.03] If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 3x + y = 6 4x − 5y = 10 Multiply the first equation by 5. Multiply the first equation by −5. Multiply the first equation by −4. Add the two equations together. Question 5 (Multiple Choice Worth 4 points) [6.02] For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 −x − y = −7 −y + 7 −y − 7 y + 7 y − 7

OpenStudy (anonymous):

6.03) You would multiply the first equation by 5. This is because: (1) 3x + y = 6 (Multiply by 5 = (2) ) + 4x − 5y = 10 (2) 15x +5y = 30 + 4x - 5y = 10 (add x's, y's and 30 + 40 = (3) ) (3) 19x +0y = 40 (times y by 0 = (4) ) (4) 19x =40

OpenStudy (anonymous):

[6.02] -y +7 This is because: (1) −x − y = −7 (leave x on its own and move -y to the other side to get (2) ) (2) -x = -7 + y (multply (-7+y) by negative (-) to get (3) ) (3) x = 7 - y (switch 7 and y around to get the answer that is required in choices (4) ) (4) x = -y + 7 (answer)

OpenStudy (anonymous):

The answer would be A) Multiply the first equation by 5. because (3x + y = 6)5 4x − 5y = 10 15x + 5y = 30 + 4x - 5y = 10 19x = 40 (19x)/19 = (10)/19 x\[\approx0.53\]

OpenStudy (anonymous):

thank you and whats the foil method?? and how would you apply it